How many Americans should be surveyed in order to be 98 percent confident that the sample mean will not differ from the population mean by more than​ 3%
Answer #1

Solution,

Given that,

\hat p = 1 - \hat p = 0.5

margin of error = E = 0.03

At 98% confidence level

\alpha = 1 - 98%

\alpha = 1 - 0.98 =0.02

\alpha/2 = 0.01

Z\alpha/2 = Z0.01 = 2.326

sample size = n = (Z\alpha / 2 / E )2 * \hat p * (1 - \hat p )

= (2.326 / 0.03)2 * 0.5 * 0.5

= 1502.85

sample size = n = 1503